OpenStudy (anonymous):

I have 5 marbles numbered 1 through 5 in a bag. Suppose I take out two different balls at random. What is the expected value of the sum of the numbers on the marbles?

OpenStudy (anonymous):

it is the some of what you get times the probability you get it

OpenStudy (anonymous):

But since it asks for the sum, how would I do that part?

OpenStudy (anonymous):

you can add i believe lets try it

OpenStudy (anonymous):

the expected value of one pick is \[\frac{1+2+3+4+5}{5}=3\]

OpenStudy (anonymous):

so my guess is the expected value of two picks is 6

OpenStudy (anonymous):

How would you find that though?

OpenStudy (anonymous):

we could also work it out the long way, probably get 6 that way too

OpenStudy (anonymous):

put \(X\) as the sum of the picks then find the possible values of \(X\) could be \(X=2,X=3,X=4,X=5,X=6,X=7,X=8,X=9,X=10,\)

OpenStudy (anonymous):

find the probability of each, then multiply and add them up

OpenStudy (anonymous):

oh that is wrong sorry, you can't get some of these lets be more careful

OpenStudy (anonymous):

you can't get 2 and you can't get 10

OpenStudy (anonymous):

you can get a total of 3, that probability is \(\frac{1}{10}\)

OpenStudy (anonymous):

you can get 4, that probability is also \(\frac{1}{10}\)

OpenStudy (anonymous):

you can get 5 and that probability is \(\frac{2}{10}\)

OpenStudy (anonymous):

is it clear where i am getting these numbers from?

OpenStudy (anonymous):

sorry, just a sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I'll just message you when I'm back

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

how far did you get ?

OpenStudy (agent0smith):

"it is the some of what you get times the probability you get it " lol "the some" :D

OpenStudy (anonymous):

Let me find my notebooks, just a sec

OpenStudy (anonymous):

the some did i write that?

OpenStudy (anonymous):

oh yeah! lol

OpenStudy (agent0smith):

I gave you a medal cos I'm cracking up haha.

OpenStudy (anonymous):

it is "some of what you get" sometimes

OpenStudy (anonymous):

Annd I've managed to lose most of the work i've done.... smh

OpenStudy (anonymous):

Also, I'm wondering, in the probabilities, is something like 1,3 different from 3,1?

OpenStudy (anonymous):

ok if you want the expected value the first step is to write down a random variable, because you have to take the expected value of something

OpenStudy (anonymous):

in this case you want the expected value of the sum so put \(x=sum\)

OpenStudy (anonymous):

okay,

OpenStudy (anonymous):

the put what the possible values of \(x\) can be in this case \(x\) could be \(\{3,4,5,6,7,8,9\}\)

OpenStudy (agent0smith):

1,3 and 3,1 means there's 2 diff. ways of getting 4, which increases the probability of getting 4. But you shouldn't have to worry about that, cos that applies to all combos (eg 1,5 and 5,1, 3,2 and 2,3 etc(

OpenStudy (anonymous):

the you have to figure out what the associated probabilities are

OpenStudy (anonymous):

5 choose 2 is 10, so we can use that as the denominator for all of these that way we don't have to worry about 14 and41 for example

OpenStudy (anonymous):

Oh, okay, that makes more sense now. So then the numerator would just be the number of distinct pairs that add to each sum.

OpenStudy (anonymous):

\(P(x=3)=\frac{1}{10}\) there is only 1 way to get a 3, namely a 1 and a 2

OpenStudy (anonymous):

yes, if you use 10 as your denominator, you only need to count how many ways, not order

OpenStudy (anonymous):

1 way to get a 4: 1 3 2 ways to get 5: 1,4 or 2, 3 2 ways to get 6: 1, 5 or 2, 4

OpenStudy (anonymous):

etc

OpenStudy (anonymous):

i think it is 1, 1, 2, 2, 2, 1, 1 respectively for 3, 4, 5, 6, 7, 8, 9 notice that they add up to 10 which is good

OpenStudy (anonymous):

For 6, I got 3,3 1,5 and 2,4

OpenStudy (anonymous):

final job is to multiply and add (find the SUM)

OpenStudy (anonymous):

think carefully here

OpenStudy (anonymous):

Ohh, I forgot, no replacement.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

I got... 6

OpenStudy (anonymous):

Is it 6? I might have messed up in my calculations.

OpenStudy (anonymous):

probably is 6 because we got 6 without thinking at the beginning

OpenStudy (anonymous):

yeah it is six the quick way to do it is put \(x\) as the number chosen first, the \(E(x)=3\) and \(y\) as the second number chose so \((Ey=3\) then the expected value of the sum is \(E(x=y)=Ex+Ey=6\)

OpenStudy (anonymous):

How do you know the expected value of the second number chosen is 3?

OpenStudy (agent0smith):

Probably because it's equally probable that you get a 5, leaving 1,2,3,4, or get a 1, leaving 2,3,4,5... all those possibilities are equal and likely lead to the expectation of the second ball being 3.

OpenStudy (anonymous):

because it is symmetric

OpenStudy (anonymous):

first second second first you cannot tell them apart

OpenStudy (agent0smith):

ie it's equally as likely to have 1,3,4,5 left as 1,2,3,5 etc.

OpenStudy (agent0smith):

(after picking the first ball)

OpenStudy (anonymous):

plus, there is no real "first second" anyway the problem says "pick two" you can put two hands in and pick if the bag is big enough

OpenStudy (anonymous):

like the difference between tossing two coins once or tossing one coin twice

OpenStudy (anonymous):

Ah, I see. Also, 6 is the final answer, correct? Or am I missing a step.

OpenStudy (agent0smith):

6 is correct. in fact the expectation will be 3 times however many balls you pick out. If you pick out all 5, the sum is 15... 5*3.

OpenStudy (anonymous):

I'ma just give the medal to agent0smith... considering 100 is pretty much as high as you can go.

OpenStudy (anonymous):

But thanks to both of you guys.

OpenStudy (anonymous):

just out of curiosity how can you pick 15 balls out of a bag that contains only 5 balls?

OpenStudy (anonymous):

i tried that once, taking $15 dollar bills out of a wallet that only contained 5, didn't work though

OpenStudy (anonymous):

ooh i see, pick 5 out of 5 must be seeing things....

OpenStudy (agent0smith):

Hmm, it appears that the expectation is the same regardless of whether it's with replacement or not... the expectation on every ball is 3, it's just that with replacement you can pick over 5 balls.

OpenStudy (agent0smith):

lol "the some" if you pick 5 balls is 15, @satellite73 ;)